OpenStudy (anonymous):

HELP!!!!! 54=7^(x)+5 find for x I am not sure who to do the log base 7 in a calculator so please help!

OpenStudy (anonymous):

54-5=7^x. do you agree with me?

OpenStudy (anonymous):

YES I DO!

OpenStudy (anonymous):

49 =7^x

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

left side 49 = 7^2 is it right?. so, we have 7^2 = 7^x ---> x =2. no log at all. hihi hihi

OpenStudy (anonymous):

do you know what i mean?

OpenStudy (anonymous):

\[7^2=7^x---> x =2\]

OpenStudy (anonymous):

why would x =2??

OpenStudy (anonymous):

okay lol just got it

OpenStudy (anonymous):

good.

OpenStudy (anonymous):

thank you super MUCH!!!!

OpenStudy (anonymous):

hihihi... glad to hear that